How many conduction electrons are there in a 1.00 mm diameter gold wire that is 40.0 cm long?
Answer #1

Radius of the wire, R = 0.5 mm = 5 * 10-4 m

Area of the coil, A = \pi R2 = 3.14 * (5 * 10-4)2 m2 = 7.85 * 10-7 m2

Volume of the coil, V = A * l = (7.85 * 10-7 m2) * (0.4 m) = 3.14 * 10-7 m3

Density of the gold coil, \rho = 19.3 * 106 g m-3

Mass = Density * Volume

m = (19.3 * 106 g m-3 ) * (3.14 * 10-7 m3 ) = 6.0602 g

m = 6.0602 g

Now atomic mass of gold, A = 196.97 g mol-1

Number of moles,n = (m/A) = (6.0602 / 196.97) mol = 0.0307671219 mol

n = 0.0307671219 mol

Avogadro's number = NA = 6.022140857 × 1023 electrons

Number of electrons = n * NA = 0.0307671219 * (6.022140857 × 1023 )

N = 1.853 * 1022 electrons